# Hyperbolic Fuzzy $C$-Means with Adaptive Weight-based Filtering for Clustering in Non-Euclidean Spaces

### v1

@@ Abstract
Clustering algorithms play a pivotal role in unsupervised learning by identifying and grouping similar objects based on shared characteristics. While traditional clustering techniques, such as hard and fuzzy center-based clustering, have been widely used, they struggle with complex, high-dimensional, and non-Euclidean datasets. In particular, the Fuzzy $C$-Means (FCM) algorithm, despite its efficiency and popularity, exhibits notable limitations in non-Euclidean spaces. Euclidean spaces assume linear separability and uniform distance scaling, limiting their effectiveness in capturing complex, hierarchical, or non-Euclidean structures in fuzzy clustering. To overcome these challenges, we introduce Filtration-based Hyperbolic Fuzzy $C$-Means (HypeFCM), a novel clustering algorithm tailored for better representation of data relationships in non-Euclidean spaces. HypeFCM integrates the principles of fuzzy clustering with hyperbolic geometry and employs a weight-based filtering mechanism to improve performance. The algorithm initializes weights using a Dirichlet distribution and iteratively refines cluster centroids and membership assignments based on a hyperbolic metric in the Poincaré Disc model. Extensive experimental evaluations demonstrate that HypeFCM significantly outperforms conventional fuzzy clustering methods in non-Euclidean settings, underscoring its robustness and effectiveness.
@@
Clustering, Hyperbolic Geometry, Fuzzy $C$-means (FCM), Poincaré Disc.
1 Introduction.
@@
FCM is widely used for its efficiency and simplicity, yet it struggles with complex, high-dimensional, and non-Euclidean datasets. To mitigate these limitations, several variants have been introduced, incorporating improved objective functions and constraints, such as adaptive FCM [19], generalized FCM [20], fuzzy weighted $c$-means [21], and generalized FCM with improved fuzzy partitioning [22]. Kernel-based approaches like kernel FCM (KFCM) [5] and constrained models, including agglomerative fuzzy $k$-means (AFKM) [23], robust self-sparse fuzzy clustering (RSSFCA) [18], robust and sparse fuzzy $k$-means (RSFKM) [24], possibilistic FCM (PFCM) [25], and principal component analysis-embedded FCM (P_SFCM) [26] as well as hyperbolic extensions such as hyperbolic smoothing-based fuzzy clustering (HSFC) [27] and Integration of hyperbolic tangent and Gaussian kernels for FCM (HGFCM) [28], have also been explored. However, despite these modifications, most of these methods remain fundamentally limited in non-Euclidean spaces, as they can still partially rely on Euclidean assumptions that fail to capture the inherent geometric complexity and hierarchical structure of such data. Consequently, these approaches often exhibit suboptimal clustering performance, reinforcing the need for a more robust, geometry-aware solution.
@@
To overcome such challenges, we propose Filtration-based Hyperbolic Fuzzy $\mathbf{C}$-Means or HypeFCM, an efficient fuzzy clustering algorithm for non-Euclidean spaces. Our approach integrates fuzzy clustering principles with hyperbolic geometry and employs a weight-based filtering mechanism to improve cluster assignments. The algorithm initializes weights using a Dirichlet distribution, iteratively updates centroids and membership weights, and computes distances using the metric of the Poincaré Disc model. The Poincaré Disc metric naturally preserves hierarchical structures due to the exponential distance scaling and the negative curvature of hyperbolic space. In contrast to Euclidean space, where distances grow linearly, hyperbolic distances expand exponentially as points move away from the origin. This property allows hyperbolic space to efficiently embed tree-like and hierarchical structures, as the distances between levels of a hierarchy are naturally stretched, ensuring clear separability between different clusters. Moreover, the Poincaré Disc model maintains conformal mapping, meaning local angles and relative positioning are preserved, which helps retain structural integrity when clustering complex, high-dimensional data. These characteristics make the Poincaré Disc metric particularly well-suited for capturing latent relationships and hierarchical patterns in non-Euclidean datasets where traditional Euclidean methods fall short. A filtration step refines connections by focusing on relevant associations, followed by weight updates until convergence. The algorithm culminates in a defuzzification process, where definitive cluster assignments are determined by optimizing the final weight matrix, effectively mapping the continuous membership values to discrete cluster labels. Our approach leverages hyperbolic space properties to improve clustering performance in complex, high-dimensional datasets.
Contributions. The key contributions are:
@@
    Our proposed algorithm HypeFCM incorporates the metric of the Poincaré Disc model by embedding the datasets inside the Poincaré Disc for more efficient performance in non-Euclidean spaces.
@@
    We present a thorough convergence analysis and discuss the computational complexity analysis of the proposed method. Other necessary proofs and derivations are provided in the supplementary document.
The paper is structured as follows: Section 2 introduces the mathematical preliminaries, Section 3 reviews related methods, and Section 4 presents the proposed HypeFCM algorithm in detail. Section 5 analyzes its computational complexity, while Section 6 discusses convergence properties. Section 7 showcases experimental results on real and synthetic datasets, Section 8 contains detailed discussions of the proposed method, and finally, Section 9 concludes the paper.
2 Preliminaries.
@@
2.1 Smooth Manifolds.
2.1.1 Manifold.
A topological space $\mathcal{M}$ is locally Euclidean of dimension $d$, if for every $\sigma\in\mathcal{M}$, there exists an open neighborhood $(X_{\sigma})$ and a map $\psi_{\sigma}:X_{\sigma}\rightarrow\psi_{\sigma}(X_{\sigma})$, where $\psi_{\sigma}(X_{\sigma})\subset\mathbb{R}^{d}$ is a homeomorphism. An $d$-dimensional Manifold is Hausdorff, $2^{nd}$ countable and locally Euclidean of dimension $d$.
  • Each $(X_{\sigma},\psi_{\sigma})$ is called a chart.
  • The family $\{(X_{\sigma},\psi_{\sigma})\}_{\sigma=1}^{k}$ is an atlas.
2.1.2 Smooth Manifold.
A smooth or $\mathbb{C}^{\infty}$ manifold is a topological manifold which comes with a family $(X_{\sigma},\psi_{\sigma})$ with the sets $X_{\sigma}\subset\mathcal{M}$ and the homeomorphisms $\psi_{\sigma}:X_{\sigma}\rightarrow\mathbb{R}^{d}$ such that $\mathcal{M}=\cup_{\sigma}X_{\sigma}$ given $(X_{\sigma}\cap X_{\beta})$ is non-null and the map $(\psi_{\beta}\circ\psi_{\sigma}^{-1})$ is smooth.
2.2 Riemannian Manifold
A Riemannian manifold $(\mathcal{M},g)$, consists of a smooth $d$-dimensional manifold $\mathcal{M}$ equipped with a Riemannian metric $g$ on $\mathcal{M}$.
For each point $p\in\mathcal{M}$, $g$ assigns a bilinear form $g_{p}:T_{p}(\mathcal{M})\times T_{p}(\mathcal{M})\rightarrow\mathbb{R}$, where $T_{p}(\mathcal{M})$ is the tangent space at $p$, such that:
  • $g_{p}$ is symmetric: $g_{p}(A,B)=g_{p}(B,A)$ for all $A,B\in T_{p}(\mathcal{M})$.
  • $g_{p}$ is positive-definite: $g_{p}(A,A)>0$ for all non-zero $A\in T_{p}(\mathcal{M})$.
If $(x^{1},x^{2},\ldots,x^{d}):U\subset\mathcal{M}\rightarrow\mathbb{R}^{d}$ are smooth local coordinates on $\mathcal{M}$, the vectors $\left\{\left(\frac{\partial}{\partial x^{1}}\right)_{p},\left(\frac{\partial}{\partial x^{2}}\right)_{p},\ldots,\left(\frac{\partial}{\partial x^{d}}\right)_{p}\right\}$ form a basis of the vector space $T_{p}(\mathcal{M})$ for any $p\in U$.
2.2.1 Length and Distance
[…]
Let $(\mathcal{M},g)$ be a connected Riemannian Manifold. Given two points $\mathbf{a,b}\in\mathcal{M}$, the distance between $\mathbf{a,b}$ is defined as: $d_{g}(\mathbf{a,b})=inf\{L(\gamma):\gamma$ is a piecewise smooth curve with $\gamma(0)=\mathbf{a},\gamma(1)=\mathbf{b}\}$. This distance is called the geodesic distance.
Let $\varsigma(\mathcal{M})$ denote the space of all $\mathbb{C}^{\infty}$ vector fields on $\mathcal{M}$. Let $\nabla:\varsigma(\mathcal{M})\times\varsigma(\mathcal{M})\rightarrow\varsigma(\mathcal{M})$ be a bi-linear map, therefore the following must hold,
  • $\nabla_{(lX_{1}+l^{\prime}X_{2})}X_{3}=l\nabla_{X_{1}}X_{3}+l^{\prime}\nabla_{X_{2}}X_{3},\forall X_{1},X_{2},X_{3}\in\varsigma(\mathcal{M})$ and $l,l^{\prime}\in\mathbb{C}^{\infty}(\mathcal{M}).$
  • $\nabla_{X_{1}}lX_{2}=X_{1}(l)X_{2}+l\nabla_{X_{1}}X_{2},\forall l\in\mathbb{C}^{\infty}(\mathcal{M}).$
$\nabla$ is said to preserve the metric and torsion-free if it satisfies the following,
  • $X_{1}(g(X_{2},X_{3}))=g(\nabla_{X_{1}}X_{2},X_{3})+g(X_{2},\nabla_{X_{1}}X_{3})$.
  • $\nabla_{X_{1}}X_{2}-\nabla_{X_{2}}X_{1}=[X_{1},X_{2}]$.
Then, $\nabla$ on $(\mathcal{M},g)$ is called the Levi-Civita connection.
2.2.2 Curvature Operator
The Riemannian curvature $\mathcal{R}$ on $\mathcal{M}$ is a trilinear map defined by: $\mathcal{R}((X_{1},X_{2}),X_{3})=\nabla_{X_{1}}\nabla_{X_{2}}X_{3}-\nabla_{X_{2}}\nabla_{X_{1}}X_{3}-\nabla_{[X_{1},X_{2}]}X_{3}\quad\forall X_{1},X_{2},X_{3}\in\varsigma(\mathcal{M}).$
2.2.3 Sectional Curvature
Let $\pi=(p,\mathscr{X})$, where $p\in\mathcal{M}$ and $\mathscr{X}\subset T_{p}(\mathcal{M})$ be a two-dimensional subspace spanned by ${A_{1},A_{2}}$. The sectional curvature of $\mathcal{M}$ is defined as,
$\mathbf{k}{(\pi)}=\frac{\langle\mathcal{R}(A_{1},A_{2})A_{2},A_{1}\rangle}{\langle A_{1},A_{1}\rangle\langle A_{2},A_{2}\rangle-\langle A_{1},A_{2}\rangle}.$
@@ 2.3 Hyperbolic space
A Hyperbolic space, represented as $\mathbb{H}^{n}$, is a non-Euclidean space of dimension $n$ which is characterized as a simply-connected Riemannian manifold with a constant negative sectional curvature $-1$. The Killing-Hopf Theorem [31] affirms that any two such Riemannian manifolds are isometrically equivalent. We will briefly discuss four of them here.
  1. 1.
    Poincaré Half Space Model: This is the upper half plane $\mathbb{H}^{n}:=\left\{\left(x_{1},x_{2},\ldots,x_{n}\right)\mid x_{n}>0,x_{i}\in\right.$ $\mathbb{R},\forall i\in\{1,2,\ldots,n\}\}$. For $\mathbf{X}:=\left(x_{1},x_{2},\ldots,x_{n}\right)$ and $\mathbf{Y}:=\left(y_{1},y_{2},\ldots,y_{n}\right)$. $\mathbf{\tilde{X}}:=\left(x_{1},x_{2},\ldots,x_{n-1},-x_{n}\right)$, the reflection of $\mathbf{X}$ with respect to the plane $x_{n}=0$ then, the metric on this space is defined as follows:
    $d\left(\mathbf{X},\mathbf{Y}\right)$ $:=2\sinh^{-1}\frac{\left\|\mathbf{Y}-\mathbf{X}\right\|}{2\sqrt{x_{n}y_{n}}}$ (1)
    $=2\log\frac{\left\|\mathbf{Y}-\mathbf{X}\right\|+\left\|\mathbf{Y}-\mathbf{\tilde{X}}\right\|}{2\sqrt{x_{n}y_{n}}}.$
  2. 2.
@@
    $d\left(\mathbf{X,Y}\right):=\cosh^{-1}\left(1+\frac{2\left\|\mathbf{X}-\mathbf{Y}\right\|^{2}}{\left(1-\left\|\mathbf{X}\right\|^{2}\right)\left(1-\left\|\mathbf{Y}\right\|^{2}\right)}\right).$ (2)
  3. 3.
    Beltrami-Klein Model: This is a model hyperbolic space in which points are represented by the points in the open unit disc in $\mathbb{R}^{n}$ and lines are represented by chords, a special type of straight lines with ideal endpoints a, b on the unit sphere. For $\mathbf{x,y}\in\mathbb{D}^{n}$, the open unit ball, the distance in this model is given as
    $d(\mathbf{x},\mathbf{y}):=\frac{1}{2}\log\left(\frac{\|\mathbf{a}\mathbf{y}\|\|\mathbf{x}\mathbf{b}\|}{\|\mathbf{a}\mathbf{x}\|\|\mathbf{y}\mathbf{b}\|}\right),$ (3)
    where $\mathbf{a,b}$ are ideal points on the boundary of the sphere of the line $(\mathbf{x,y})$ with $\|\mathbf{a}\mathbf{y}\|>\|\mathbf{a}\mathbf{x}\|$ and $\|\mathbf{x}\mathbf{b}\|>\|\mathbf{b}\mathbf{y}\|$.
  4. 4.
    Hyperboloid Model: This is a model of hyperbolic space also known as the Minkowski Model, which is the forward sheet $S^{+}$of the two hyperbolic sheets embedded in the $(n+1)$ dimensional Minkowski Space. The hyperbolic distance between two points $\mathbf{X}=\left(x_{0},x_{1},\ldots,x_{n}\right)$ and $\mathbf{Y}=\left(y_{0},y_{1},\ldots,y_{n}\right)$ is given as,
    $d(\mathbf{X,Y}):=\cosh^{-1}(-M(\mathbf{X,Y})).$ (4)
    where,
    $M\left(\left(x_{0},x_{1},\ldots,x_{n}\right),\left(y_{0},y_{1},\ldots,y_{n}\right)\right)$
    $\qquad\qquad=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}.$
    Here, $M$ is simply the Minkowski Dot Product.
2.4 Gyrovector spaces.
The framework of gyrovector spaces establishes an elegant non-associative algebraic structure that naturally captures and formalizes the properties of hyperbolic geometry analogous to the vector spaces provide the algebraic setting for Euclidean geometry [32]. We denote $\mathbb{D}_{\alpha}^{p}:=\left\{\mathbf{v}\in\mathbb{R}^{p}\mid\alpha\|\mathbf{v}\|^{2}<1\right\}$ taking $\alpha\geq 0$. If $\alpha=0$, then $\mathbb{D}_{\alpha}^{p}=\mathbb{R}^{p}$; if $\alpha>0$, then $\mathbb{D}_{\alpha}^{p}$ is the open ball of radius $1/\sqrt{\alpha}$. If $\alpha=1$ then we recover the unit ball $\mathbb{D}^{p}$.
The operations in Poincaré Disc Model are well-defined, computationally efficient, as most of the geometric operations have simple closed-form expressions in Poincaré Disc Model, and it also preserves the angle. In this context, we will provide a brief discussion on Möbius gyrovector addition and Möbius scalar multiplication on the Poincaré Disc model. Due to the invariance of geometric properties under isometric mappings between the hyperbolic spaces, the fundamental additive and multiplicative algebraic structures can be isomorphically translated across different models of hyperbolic geometry while preserving their essential characteristics [33]. We have established the isometric equivalence between the Poincaré Disc model and the hyperboloid model by explicitly constructing a mapping between them. Let $\mathbb{D}^{n}=\{x\in\mathbb{R}^{n}:\|x\|<1\}$ be the Poincaré Disc model, and let the hyperboloid model be
@@
is an isometry. It preserves the hyperbolic distance and induces the same Riemannian metric. Hence, we will be using the Poincaré Disc model throughout.
@@
Möbius scalar multiplication. For $\alpha>0$, the Möbius scalar multiplication of $\mathbf{v}\in\mathbb{D}_{\alpha}^{p}\backslash\{\boldsymbol{0}\}$ by a real number $\lambda\in\mathbb{R}$ is defined as,
@@
$d_{H}(\mathbf{v},\mathbf{w})=(2/\sqrt{\alpha})\tanh^{-1}\left(\sqrt{\alpha}\left\|-\mathbf{v}\oplus_{\alpha}\mathbf{w}\right\|\right).$ (7)
@@
where ${\lambda_{\mathbf{x}}}=\frac{2}{1-\lambda{\|x\|}^{2}}$ and $\bigoplus_{\alpha}$ represents the Möbius gyrovector addition as Equation 5, $n$ is the number of data points, $m$ is the fuzziness parameter.
@@
As the parameter $\alpha$ approaches to 00,
@@
3 related works.
3.1 Fuzzy $C$-means.

### v2

# Hyperbolic Fuzzy C-Means with Adaptive Weight-based Filtering for Efficient Clustering

@@ Abstract
Clustering algorithms play a pivotal role in unsupervised learning by identifying and grouping similar objects based on shared characteristics. Although traditional clustering techniques, such as hard and fuzzy center-based clustering, have been widely used, they struggle with complex, high-dimensional, and non-Euclidean datasets. In particular, the fuzzy $C$-Means (FCM) algorithm, despite its efficiency and popularity, exhibits notable limitations in non-Euclidean spaces. Euclidean spaces assume linear separability and uniform distance scaling, limiting their effectiveness in capturing complex, hierarchical, or non-Euclidean structures in fuzzy clustering. To overcome these challenges, we introduce Filtration-based Hyperbolic Fuzzy $C$-Means (HypeFCM), a novel clustering algorithm tailored for better representation of data relationships in non-Euclidean spaces. HypeFCM integrates the principles of fuzzy clustering with hyperbolic geometry and employs a weight-based filtering mechanism to improve performance. The algorithm initializes weights using a Dirichlet distribution and iteratively refines cluster centroids and membership assignments based on a hyperbolic metric in the Poincaré Disc model. Extensive experimental evaluations on $6$ synthetic and $12$ real-world datasets demonstrate that HypeFCM significantly outperforms conventional fuzzy clustering methods in non-Euclidean settings, underscoring its robustness and effectiveness.
@@
Clustering, Hyperbolic Geometry, Fuzzy $C$-means (FCM), Poincaré Disc, Filtration.
1 Introduction
@@
FCM is widely used for its efficiency and simplicity, yet it struggles with complex, high-dimensional, and non-Euclidean datasets. To mitigate these limitations, several variants have been introduced, incorporating improved objective functions and constraints, such as adaptive FCM [19], generalized FCM [20], fuzzy weighted $c$-means [21], and generalized FCM with improved fuzzy partitioning [22]. Kernel-based approaches like kernel FCM (KFCM) [5] and constrained models, including agglomerative fuzzy $k$-means (AFKM) [23], robust self-sparse fuzzy clustering (RSSFCA) [18], robust and sparse fuzzy $k$-means (RSFKM) [24], possibilistic FCM (PFCM) [25], principal component analysis-embedded FCM (P_SFCM) [26] as well as hyperbolic extensions such as hyperbolic smoothing-based fuzzy clustering (HSFC) [27], Integration of hyperbolic tangent and Gaussian kernels for FCM (HGFCM) [28], have also been explored. However, despite these modifications, most of these methods remain fundamentally limited in non-Euclidean spaces, as they can still partially rely on Euclidean assumptions that fail to capture the inherent geometric complexity and hierarchical structure of such data. Consequently, these approaches often exhibit suboptimal clustering performance, reinforcing the need for a more robust, geometry-aware solution. Recently, Wu and Pan proposed FCPFM [29], a fuzzy clustering algorithm in hyperbolic space that utilizes Fréchet mean for centroid updates, and introduced a kernelized variant of it.
@@
In this paper, we propose Filtration-based Hyperbolic Fuzzy $\mathbf{C}$-Means or HypeFCM, a novel fuzzy clustering algorithm for non-Euclidean spaces. Our approach not only integrates fuzzy clustering principles with hyperbolic geometry but also employs a weight-based filtering mechanism to improve cluster assignments. A defining strength of our algorithm lies in ensuring that the centroid update remains inherently constrained within the geometry of the Poincaré ball, thereby preserving the integrity of hyperbolic space throughout the clustering process. The algorithm initializes weights using a Dirichlet distribution, iteratively updates centroids and membership weights, and computes distances using the metric of the Poincaré Disc model. The Poincaré Disc metric naturally preserves hierarchical structures due to the exponential distance scaling and the negative curvature of hyperbolic space. In contrast to Euclidean space, where distances grow linearly, hyperbolic distances expand exponentially as points move away from the origin. This property allows hyperbolic space to efficiently embed tree-like and hierarchical structures, as the distances between levels of a hierarchy are naturally stretched, ensuring clear separability between different clusters. Moreover, the Poincaré Disc model maintains conformal mapping, meaning local angles and relative positioning are preserved, which helps retain structural integrity when clustering complex, high-dimensional data. These characteristics make the Poincaré Disc metric particularly well-suited for capturing latent relationships and hierarchical patterns in non-Euclidean datasets where traditional Euclidean methods fall short. A filtration step refines connections by focusing on relevant associations, followed by weight updates until convergence. The algorithm culminates in a defuzzification process, where definitive cluster assignments are determined by optimizing the final weight matrix, effectively mapping the continuous membership values to discrete cluster labels.
Contributions. Our method enjoys the following advantages:
@@
    Our proposed algorithm HypeFCM incorporates the metric of the Poincaré Disc model by embedding the datasets inside the Poincaré Disc and ensuring the subsequent centroid updates are also constrained within the geometry of the Poincaré ball.
@@
    We present a thorough convergence analysis and discuss the computational complexity analysis of the proposed method. Other necessary proofs and derivations are provided in the Appendix.
Table I: Notations.
Notation Description
$\mathbf{X}\in\mathbb{R}^{n\times p}$ The given dataset
$\mathbf{V}\in\mathbb{R}^{c\times p}$ The cluster centroid matrix
$\mathbf{W}\in\mathbb{R}^{n\times c}$ The membership matrix
$\mathbf{x}_{i}\in\mathbb{R}^{p}$ The $i$-th sample in $\mathbf{X}$
$\mathbf{v}_{j}\in\mathbb{R}^{p}$ The $j$-th cluster centroid in $\mathbf{V}$
$\mathbf{1}_{c}$ All-ones vector with a length of $c$
$n$ Number of samples in $\mathbf{X}$
$p$ Dimensionality of $\mathbf{X}$
$c$ Number of clusters in $\mathbf{X}$
$m$ The value of the fuzziness parameter
$T$ Maximum number of iterations
$k$ The Filtration value in the algorithm
$\operatorname{Tr}(\cdot)$ Denotes the trace of a matrix
$\alpha$ The curvature parameter of the Poincaré Disc model
$\varepsilon$ The tolerance value
The paper is structured as follows: Section 2 introduces the mathematical preliminaries, Section 3 reviews related works, Section 4 presents the proposed HypeFCM algorithm in detail, Section 5 analyzes its computational complexity, while Section 6 discusses convergence analysis of our method. Section 7 presents experimental results on real and synthetic datasets. Section 8 provides detailed discussions of the proposed method, and Section 9 concludes the paper.
2 Preliminaries
@@ 2.1 Hyperbolic space
A Hyperbolic space, represented as $\mathbb{H}^{n}$, is a non-Euclidean space of dimension $n$ which is characterized as a simply-connected Riemannian manifold with a constant negative sectional curvature 1-1. The Killing-Hopf theorem [32] affirms that any two such Riemannian manifolds are isometrically equivalent. We will briefly discuss the Poincaré Disc Model here.
@@
$d\left(\mathbf{X,Y}\right)=\cosh^{-1}\left(1+\frac{2\left\|\mathbf{X}-\mathbf{Y}\right\|^{2}}{\left(1-\left\|\mathbf{X}\right\|^{2}\right)\left(1-\left\|\mathbf{Y}\right\|^{2}\right)}\right).$ (1)
2.2 Gyrovector Spaces
The framework of gyrovector spaces establishes an elegant non-associative algebraic structure that naturally captures and formalizes the properties of hyperbolic geometry, analogous to the way vector spaces provide the algebraic setting for Euclidean geometry [33]. We denote $\mathbb{D}_{\alpha}^{p}:=\left\{\mathbf{v}\in\mathbb{R}^{p}\mid\alpha\|\mathbf{v}\|^{2}<1\right\}$ taking $\alpha\geq 0$. If $\alpha=0$, then $\mathbb{D}_{\alpha}^{p}=\mathbb{R}^{p}$; if $\alpha>0$, then $\mathbb{D}_{\alpha}^{p}$ is the open ball of radius $1/\sqrt{\alpha}$. If $\alpha=1$ then we recover the unit ball $\mathbb{D}^{p}$.
The operations in Poincaré Disc Model are well-defined, computationally efficient as most of the geometric operations have simple closed-form expressions in Poincaré Disc Model. In this context, we will provide a brief discussion on möbius gyrovector addition and möbius scalar multiplication on the Poincaré Disc model. Due to the invariance of geometric properties under isometric mappings between the hyperbolic spaces, the fundamental additive and multiplicative algebraic structures can be isomorphically translated across different models of hyperbolic geometry while preserving their essential characteristics [34]. We have established the isometric equivalence between the Poincaré Disc model and the hyperboloid model by explicitly constructing a mapping between them. Let $\mathbb{D}^{n}=\{x\in\mathbb{R}^{n}:\|x\|<1\}$ be the Poincaré Disc model, and let the hyperboloid model be,
@@ 2.2 Gyrovector Spaces
is an isometry. It preserves the hyperbolic distance and induces the same Riemannian metric. Hence, we will be using the Poincaré Disc model throughout [35] [36].
@@
[…]
@@
$d_{hyp}(\mathbf{v},\mathbf{w})=(2/\sqrt{\alpha})\tanh^{-1}\left(\sqrt{\alpha}\left\|-\mathbf{v}\oplus_{\alpha}\mathbf{w}\right\|\right).$ (4)
@@
where ${\lambda_{\mathbf{x}}}=\frac{2}{1-\lambda{\|x\|}^{2}}$ and $\bigoplus_{\alpha}$ represents the Möbius gyrovector addition as Equation 2.
@@
As the parameter $\alpha$ approaches 0,
@@
3 related works
3.1 Fuzzy $C$-means (FCM)